Find the variation constant and an equation of variation for the given situation:
-y varies directly as x, and y=0.1 when x=0.2

Answer by stanbon(57361)

Find the variation constant and an equation of variation for the given situation:
-y varies directly as x, and y=0.1 when x=0.2
---
y = kx
Find "k":
0.1 = k(0.2)
k = 1/2
------
Equation:
y = (1/2)x
===================
Cheers,
Stan H.